OpenStudy (anonymous):

if f(x)=8-2x^2,how many integers satisfy the condition that f(x) is positive?

OpenStudy (hoblos):

8-2x^2 >0 2(4-x^2)>0 4-x^2 >2 x^2< 4 -2<x<2

OpenStudy (hoblos):

they are 3.. -1,0,1

OpenStudy (anonymous):

thanks.:)

OpenStudy (hoblos):

my pleasure XD

OpenStudy (campbell_st):

let f(x) = 0 then 8 - 2x^2 = 0 or 8 = 2x^2 then 4 = x^2 then \[x = \pm 2\] or \[-2\le x \le 2\]